What is the molar solubility of barium hydroxide in water? Would you characterize barium hydroxide as being soluble? Why? 3. 4. What is the molar solubility of barium hydroxide in a pH 9.0 buffer?
Answer #1

3. Molar solubility is of any compound is the number of moles that can be dissolved per liter of solution until it is saturated . It can be calculated if substance's solubility product constant (Ksp) and stoichiometry are known. The units are mol/L (written as M). Ksp is an equilibrium constant for equilibrium between a solids and its ions in a solution. The higher the value of Ksp, the more soluble the compound.

Ba(OH)2 ionises in water as  

Ba(OH)2 \rightleftharpoons Ba2+ + 2HO-

Its Ksp value is 2.55 x 10-4 at 25 oC , So Ksp = [Ba2+] [ HO-]2

According to the balanced equation, for every one mole of Ba2+ formed, 2 moles of HO- are formed. Let's assign the variable S for the solubility of the Ba2+

[Ba2+] = S [ HO-]2 = 2S So, Ksp = S x (2S)2 = 4S3 = 2.55 x 10-4

Molar solubility of Ba(OH)2 is S = 0.04 mole/L

Its molar solubility is very low. So, It can not be said as "Soluble".

4. We know

pH + pOH = 14
pOH = 14-9 =5 (given pH =9)
pOH = -log[OH-]
[OH-] = 10-pOH= 10-5.....Eq 1

Now   [Ba+2] = S and [OH-]= 2S, So molar solubility at pH = 9 is 
S = 10-5/2 = 0.5 x 10-5  mole/L